OpenStudy (anonymous):

find the exact value of ln e^2.50 - ln sqrt e?

geerky42 (geerky42):

Just remember that \(\log a^b = b~\log a\) and \(\sqrt c = c^{1/2}\)

OpenStudy (anonymous):

Thanks!

geerky42 (geerky42):

Ok, so you can figure it out now?

OpenStudy (anonymous):

Not really

geerky42 (geerky42):

OK, so \(\ln e^{2.5} - \ln e^{0.5} = 2.5\ln e - 0.5\ln e\), right? And since \(\ln e = 1\), we have \(2.5 - 0.5 = \boxed{2}\)

OpenStudy (anonymous):

Thank you!